UT 20151102 D/E night

luci2.20151102.NNNN.fits

0001 test image with blind filters 1x5.0s

Error moving mask to storage.

Peter online working on LUCI1, so he was able to fix it fast.

### LUCI2 Observing UT 20151102
luci2.20151102.NNNN.fits

0001 1x5.0s dark
test image with blind filters 1x5.0s
0002 1x3.0s dark (for transform check from prev. night)

Error moving mask to storage.
Peter online working on LUCI1,
he was able to fix it fast.

Init all on LUCI2
No restart of scheduler

12 deg twilight @

Recentered N375 field stop on detector
Initial collimation @ AM=1.85, FWHM=
Two odd jumps during collimation, star blew up to ~2"
DGH cleared all active optics and restarted

01:28 Beltran
S875-C_H2.img.xml
12x10.0s each exposure (way too long for a standard)

Guide star not found...there but defocussed again
Going back to initial correction star...

01:37 Beltran
S875-C_H2.img.xml
0003 0007 12x10.0s star has ~413k ADU peak...34k/DIT
No dither between 3 and 4? Others are
FWHM=0.55! 3.5pix on LUCI AM=2.19 @ end
G* elongated on guider, but images on LUCI look round

S875-C_Ks.img.xml
0008 0012 20x3.0s (also way too long for a standard)
looks close to saturated in 1 DIT
4.2 pixels FWHM

2:01 G35.20-0.74N (A1) A1_H2.img.xml
0013 0030
AM 1.45 @ start FWHM=0.5" on guider
~6k ADU peak in central star/blob <4pix on LUCI

02:59 G35.03+0.35 (A2) A2_H2.img.xml
0031 0045
AM 1.88 @ start FWHM=0.5
~5k ADU peak per DIT 4.7pix FWHM (~0.6")
AM 2.54 @ end FWHM=5 pix
03:45 NGC34_V2.xml

03:46 another MOS error
called James
Looks to me like initialize will clear error

04:34 James found the spindle resolver "noisy"
reading values from 129 to 132 (should be 131.3)

04:49 canceled preset
Move to zenith, ROT=210
spindle resolver still noisy

04:54 Initializing the MOS
spindle resolver still noisy
Checked LUCI1's spindle resolver and it is stable

05:12 We decided that it is not safe to continue with LUCI2 so
we are switching to LBCs.

Blind filters, N375, no mask, Mirror in the beam.

### LBCs:

NGC828_copoint.ob
- OB takes all 4 chips!

06:37 NGC828_IB_5ptL.ob
B 2.9 pixels!
R 2.5 pixels FWHM!
Data looks good

06:56 DOFPIA, /X2 @ copoint position
419 28 75 152 -44 -120 -12 0.34 0.21 red
711 -172 -186 29 -175 -163 26 0.85 0.33 blue
-98 -15 -100 -15 -20 3 1 0.33 0.09 red - converged
413 -65 -186 -94 -237 -135 0 0.86 0.24 blue
-103 0 -45 49 75 -22 -7 0.34 0.10 red
69 51 -168 -9 -65 -67 5 0.88 0.11 blue - converged

07:02 NGC828_IB_5ptR.ob
B 3.3 pix
R 2.3 pix! 0.5"
Data looks good, FWHM increased slightly through OB

07:20 SA114 EL~25 (no co-point, short exposures)
DOFPIA, /X2
9942 0 0 0 0 0 0 0.52 3.92 red
4255 -107 1 -511 167 86 -31 0.78 1.70 blue
616 292 178 134 -337 -1 -36 0.53 0.32 red
-218 -248 -33 386 -364 244 27 0.97 0.32 blue - converged
-156 139 132 -125 -21 -95 -32 0.54 0.14 red - converged
-614 -225 35 290 -3 1 -6 0.92 0.29 blue

07:31 SA114_BVRI.ob EL~22.5
Not many LBCB exposures here??

07:36 Galadi1 EL~57 (no co-point, short exposures)
DOFPIA, /X2
-3613 -249 -296 43 -118 19 68 0.32 1.43 red
-1898 170 39 139 834 -248 32 0.77 0.86 blue
-119 -129 -134 -84 13 -127 32 0.46 0.14 red - converged
-49 32 112 -188 -129 -37 38 0.97 0.14 blue - converged

07:41 Galadil-BVRI.ob
Also only LBCR exposures ??
~2.5 pixels

07:48 NGC1502 - not that far from the moon
DOFPIA, /X2 # Odd scattered light pattern in backgrounds
673 -172 235 -201 207 -224 3 0.44 0.35 red
546 15 148 252 -359 -366 -17 0.85 0.38 blue
-20 -39 53 10 26 -105 1 0.42 0.11 red - converged
468 125 -137 -18 52 -235 -34 0.78 0.26 blue
323 -96 -49 66 -84 -140 -6 0.44 0.18 red
-614 143 -104 57 -33 169 18 0.66 0.28 blue
-308 -78 125 -98 49 -1 2 0.47 0.16 red
55 -119 42 -18 -190 -81 -13 0.93 0.13 blue - converged

07:59 WT10_395 to copoint

08:02 NGC1502_9ptL.ob
B 2.8 pix
R 2.3 pix

08:13:03.788175 OB execution error: one or more systems failed [src/lbc/actions.c:1160]
08:13:03.687102 >>> preset/expose/save interrupted at preset time, retcode=7
08:12:52.586994 filter is more than 10 times the maximum allowed distance from its nominal position [src/filters/filters.c:2334]
08:12:51.286887 timeout error on motion, took 9.960211 secs [src/filters/filters.c:2246]
IT 5548 says just restart OB:
08:22 DOFPIA, /X2
533 93 7 211 -195 -60 -5 0.39 0.26 red
0 -26 -101 -36 -25 -167 -6 0.82 0.14 blue
339 54 -8 -24 77 -108 -13 0.46 0.17 red
82 203 64 -48 50 -33 13 0.77 0.11 blue - converged
169 -31 -32 -74 -6 -117 -12 0.42 0.13 red - converged
-204 -32 -110 36 181 52 8 0.75 0.15 blue

08:28 NGC1502_9ptL.ob again
B 2.6 pix 2.1! 2.5
R 2.6 pix 2.3 2.8
OB runs for well over 30 minutes (about 45).
@30 min mark 48/72 done

09:11 cvOB7 2014LA27 5:17:01.001 36:03:29.02
DOFPIA, /X2
79 100 -168 261 8 -127 0 0.38 0.18 red - converged
364 -72 183 23 79 -168 -8 0.68 0.21 blue
185 -21 -58 -100 29 -97 -2 0.38 0.13 red
239 51 151 -134 -1 -138 -14 0.69 0.17 blue - converged

09:17 cvOB7.ob AIR 1.006 @ end
B 3.0 pix
R 2.5 pix

09:27 NGC 1502
DOFPIA, /X2
-825 97 314 -96 -170 -108 0 0.43 0.37 red
-2055 124 -24 103 169 -160 -9 0.74 0.83 blue
52 37 233 -41 -77 -110 -24 0.47 0.13 red - converged
76 -6 -71 159 -97 -87 -3 0.77 0.13 blue - converged

09:33 NGC1502_9ptR.ob
B 2.2 pix
R 2.2 pix
@28/72 blue IQ looks degraded...aborted

WT10_395 Copoint
09:54 NGC1502_9ptL.ob
Rotator error:
09:54:33.432914 QueryRotator failed, trajectory is invalid
Since this OB was aborted, just continue...
DOFPIA, /X2
536 42 92 27 65 -129 -16 0.38 0.24 red
316 139 62 80 -46 -87 -7 0.72 0.17 blue
40 17 -57 506 17 -16 -3 0.39 0.24 red
-46 48 -116 0 -30 -76 -1 0.77 0.10 blue - converged
-85 44 8 -333 2 -68 -11 0.38 0.18 red - converged
120 -28 -112 59 65 -3 0 0.72 0.11 blue

10:02 NGC1502_9ptR.ob
B 2.2 pix 2.6 2.8
R 2.4 pix 2.3 2.5

10:45 cvOB7
10:45:05.613277 QueryRotator failed, trajectory is invalid
Didn't seem to slow things down much
DOFPIA, /X2
338 -142 -158 327 -35 -158 -7 0.27 0.24 red
1403 -312 95 -16 -352 -219 -5 0.90 0.60 blue
607 -77 -86 -25 107 -283 3 0.30 0.31 red
74 -28 243 -7 64 -110 -23 0.77 0.13 blue - converged
-100 -45 -24 -21 -56 -76 -17 0.32 0.10 red - converged
-117 -65 125 -5 42 -16 4 0.73 0.10 blue

10:52 cvOB7.ob AIR 1.05 @ end
B 3.3
R 3.0
11:03 cvOB10 2013EU9
DOFPIA, /X2 .. forgot to abort prev. double Sph added
618 -42 -58 78 -100 -218 0 0.27 0.30 red
628 -169 38 104 290 -356 -27 0.57 0.37 blue
496 -25 -40 97 79 -218 -1 0.26 0.26 red
-210 -183 -122 -120 -30 -85 5 0.57 0.15 blue - converged
277 -33 -4 -85 36 -172 -4 0.25 0.18 red
201 203 103 -17 7 -99 6 0.59 0.15 blue
435 -23 15 -45 17 -177 -11 0.27 0.22 red
68 66 130 -11 -38 -45 -3 0.56 0.10 blue
-56 -29 16 10 -12 -22 -10 0.27 0.08 red - converged
228 -137 -64 68 -26 -37 -3 0.53 0.13 blue

11:16 cvOB10.ob 2013EU9
B 3.0 pix
R 2.4

SA95 The specifiedimported OB specifies NO filter and then NO image

11:28 UGC 4635 no FC
11:30 WT10_348 for copointing

11:33 UGC 4635
1275 63 -42 -257 118 -308 -4 0.37 0.56 red
997 113 -102 -37 -111 -228 -42 0.72 0.43 blue
-221 47 -7 170 62 -3 0 0.39 0.14 red - converged
0 -97 -1 39 19 3 -18 0.74 0.09 blue - converged

11:38 UGC_4635_IB_5ptL.ob
B 3.1
R 2.8

11:56 cvOB10.ob 2013EU9
DOFPIA, /X2
-598 55 1 86 -345 -16 29 0.28 0.30 red
-188 152 163 -97 485 -206 16 0.52 0.29 blue
176 9 33 -56 -87 -44 -6 0.32 0.12 red - converged
-16606 0 0 0 0 0 0 0.65 6.54 blue
-152 -46 -17 165 -27 -55 6 0.28 0.13 red
-225 -23 157 -17 -94 -112 10 0.49 0.15 blue

12:06 cvOB10.ob 2013EU9
B 3.0 pix
R 2.6 pix

Astron. Twilight: 05:15 MST = UT 12:15

Twilight Flats - Field: CAblank2 (09 13 49 +46 13 58)
V-BESSEL and R-BESSEL tests
1000 500 ADU for 1 sec
1200/3sec 1500/10sec
1300/5sec 1300/8sec
Flats @ PA=0 ~1300 ADU
Flats @ PA=180 ~1500 ADU
1750/5sec 1900/8sec
2000 2300
2200 2000 5sec ea
2700 2700
Flats @ PA=180 ~3800 to ~5200 ADU
Flats @ PA= 0 ~5600 to ~9300 ADU

g-SLOAN and r-SLOAN
Flats @ PA= 0 ~8k to ~15k ADU (2 sec ea)
Flats @ PA=180 ~10k to ~21k ADU (1 sec ea)

B-BESSEL and I-BESSEL
Flats @ PA=180 ~12k to ~23k ADU (1 sec ea) B-BESSEL
Most of the I-BESSEL probably too bright (29k-51k)?
Flats @ PA= 0 ~28k to ~65k ADU (1 sec ea) B-BESSEL
Most of the I-BESSEL and last 1-2 B-BESSEL probably saturated

01:15 Done with twilight flats, closing

Somewhere in the twilight flats:

13:00:14.341000 CAMERA CCDCTRL too few downloaded pixels (42465372 instead of 42467328) [ccdctrl.c:2758]

-- %USERSIG{DavidThompson - 2015-11-02}%

Comments

 
Topic revision: r2 - 02 Nov 2015, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback